OpenStudy (anonymous):

What is the domain and range of the quadratic equation y = –x2 – 14x – 52? a. Domain: All Real Numbers Range: y –3 b. Domain: All Real Numbers Range: y –7 c. Domain: All Real Numbers Range: y –3 d. Domain: All Real Numbers Range: y –7

OpenStudy (amistre64):

well, there seems to be no dispute on domain; so what does range mean?

OpenStudy (anonymous):

equal then or greater to

OpenStudy (amistre64):

close; range is all possible outcomes; can be greter than ot less than a specific boundary; or the set of all reals as well

OpenStudy (amistre64):

in this case; id see if -3 or -7 produces a "bigger" y value and pick that bigger one

OpenStudy (amistre64):

then youll have to determine which symbol is where since youre copy paste of the question omits those parts

OpenStudy (anonymous):

The vertex would be at \[(\frac{-b}{2a},f(\frac{-b}{2a}))=(-7,-3)\]. The graph of this function is opening down, so -3 is a maximum. Range is: \[(-\infty, -3]\]